OpenStudy (anonymous):

A birthday cake is cut into 16 equal pieces. If the cake has a diameter of 14 inches, what is the area of one piece of cake?

OpenStudy (anonymous):

A = pi * r^2 A = pi * 7^2 A = 49 pi 16 equal pieces: 49 pi / 16 So around 9.621127502 inches.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Become a fan!

OpenStudy (anonymous):

k